OpenStudy (anonymous):

Prove the identity : sin (x-y)sin(x+y) = sin^2 x - sin^2 y

OpenStudy (anonymous):

@ybarrap

OpenStudy (kc_kennylau):

\[\sin(A+B)=\sin A\cos B+\cos A\sin B\\\sin(A-B)=\sin A\cos B-\cos A\sin B\]

OpenStudy (ybarrap):

use sin(a-b)=sin(a) cos(b)-cos(a) sin(b) and sin(a+b)=sin(a) cos(b)+cos(a) sin(b) $$ \left(\sin(a) \cos(b)-\cos(a) \sin(b)\right)\times \left (\sin(a) \cos(b)+\cos(a) \sin(b)\right ) $$

OpenStudy (kc_kennylau):

\[\alpha^2-\beta^2=(\alpha+\beta)(\alpha-\beta)\]

OpenStudy (ybarrap):

perfect!

OpenStudy (ybarrap):

take $$ \alpha=\sin a\cos b\\ \beta=\cos a \sin b $$

OpenStudy (anonymous):

okayy sorry I haven't gotten much to those symbols yet here

OpenStudy (anonymous):

what do they represent in this problem?

OpenStudy (kc_kennylau):

They denote that for any \(\alpha,\beta,A,B\), the formulae still hold.

OpenStudy (kc_kennylau):

For example, \(\sin^2\theta+\cos^2\theta=1\) means that for any \(\theta\) the formula holds

OpenStudy (kc_kennylau):

I know it's not proper grammar but yeah

OpenStudy (anonymous):

okayyy

OpenStudy (anonymous):

I thought I was using x and y

OpenStudy (anonymous):

thnx for the help though

OpenStudy (kc_kennylau):

okay sorry for the wrong hints, here is the answer:

OpenStudy (kc_kennylau):

\[\begin{array}{cl} &\sin^2x-\sin^2y\\ =&(\sin x+\sin y)(\sin x-\sin y)\\ =&4\sin\frac{x+y}2\cos\frac{x-y}2\cos\frac{x+y}2\sin\frac{x-y}2\\ =&\left(2\sin\frac{x+y}2\cos\frac{x+y}2\right)\left(2\sin\frac{x-y}2\cos\frac{x-y}2\right)\\ =&\sin(x+y)\sin(x-y) \end{array}\]

OpenStudy (anonymous):

awesome thank you so much! @ybarrap I don't know if you're still typing or my computer froze but thank you too :)

OpenStudy (anonymous):

I have two other questions which I'd like you guys to check out but I will write them in a different post

OpenStudy (ybarrap):

And if you want to try a longer version: $$ \alpha^2-\beta^2=(\alpha+\beta)(\alpha-\beta) $$ Is the same as $$ \left(\sin(a) \cos(b)-\cos(a) \sin(b)\right)\times \left (\sin(a) \cos(b)+\cos(a) \sin(b)\right ) \star $$ Where, $$ \alpha=\sin a\cos b\\ \beta=\cos a \sin b $$ This means that when you multiply out \(\star\) $$ \sin(a+b)\sin(a-b)\\ \left(\sin(a) \cos(b)-\cos(a) \sin(b)\right)\times \left (\sin(a) \cos(b)+\cos(a) \sin(b)\right )\\ =(\sin a\cos b)^2-(\cos a \sin b)^2\\ \sin^2a\cos^2b-\cos^2a\sin^2b\\ (1/2 (1-\cos(2 a)))1/2 (cos(2 b)+1)-1/2 (cos(2 a)+1)(1/2 (1-\cos(2 b)))\\ =(1/4)((\cos 2b+1-\cos(2a)\cos(2b)-\\ \cos(2a)-\cos(2a)+\cos(2a)\cos(2b)-1+\cos(2b))\\ =(1/4)(2\cos 2b-2\cos2a)\\ =(1/2)(\cos 2b-\cos2a)\\ =(1/2)(cos^2(b)-sin^2(b) -cos^2(a)+sin^2(a))\\ =(1/2)( 1-sin^2(b)-sin^2b -1+sin^2a+sin^2a)\\ =(1/2)(-2\sin^2b+2\sin^2a)\\ =-\sin^2b+\sin^2a\\ $$

OpenStudy (anonymous):

wow thank you both sooo much!! I think I'm understanding this better :)